The cost in dollars of manufacturing w wing nuts is given by the expression
0.07w + 13.3. Find the cost when 375 wing nuts are made. What is the average cost to manufacture on wing nut?

Answer by wuwei96815(245)

Cost = 0.07w + 13.3 I am assuming that w=number of wing nuts.

C = 0.07(375) + 13.3
C = 26.25 + 13.3
C = 39.55 I am assuming that the answer is in dollars.

Average cost = total cost/number of wing nuts
AC = 39.55/375 = 0.1054666 cents

Hope that is right.